# Decoding Chess Puzzle Play and Standard Cognitive Tasks for BCI: A Low-Cost EEG Study

### v1

@@ Abstract
While consumer-grade electroencephalography (EEG) devices show promise for Brain-Computer Interface (BCI) applications, their efficacy in detecting subtle cognitive states remains understudied. Using a combination of established cognitive paradigms (N-Back, Stroop, and Mental Rotation) and a novel ecological task (Chess puzzles), we demonstrate successful distinctions of workload levels within some tasks, as well as differentiation between task types using the MUSE 2 device. With machine learning we further show reliable predictive power to differentiate between workload levels in the N-Back task, while also achieving effective cross-task classification. These findings demonstrate that consumer-grade EEG devices can effectively detect and differentiate various forms of cognitive workload, and that they can be leveraged with some success towards real-time classification distinguishing workload in some tasks, as well as in differentiating between nuanced cognitive states, supporting their potential use in adaptive BCI applications. Research code and data are further provided for future researchers.
1 Introduction
@@
Electroencephalography (EEG) has emerged as a powerful tool in Brain-Computer Interface (BCI) research. In particular, EEG has an established capability to measure cognitive workload (Gevins and Smith, 2003), which has proven valuable in ergonomics and human-computer interaction studies for differentiating task difficulties (Zarjam et al., 2015). And recent advances in consumer-grade devices have expanded EEG’s accessibility beyond traditional clinical settings. This democratization has enabled researchers across diverse fields, from engineering to cognitive neuroscience, to leverage EEG technology in novel ways (Sabio et al., 2024).
@@
However, bridging the gap between clinical neuroscience findings, initial validation of low-cost devices, and practical applications for general users remains an active area of research (Värbu et al., 2022). Further, while EEG technology continues to advance, significant work is still needed to establish reliable real-time signal processing methods suitable for BCI applications (McFarland and Wolpaw, 2017). These challenges necessitate empirical studies to both validate the technological developments of lower-cost devices, and to inform the design of future user interfaces.
@@
Our research employs the MUSE 2, a wireless and portable EEG device initially validated for cognitive workload measurement in BCI applications (Peining et al., 2017; Bird et al., 2018). Despite this initial validation, there remains limited understanding of its practical application in realistic BCI scenarios. To address this gap, we investigate the device’s ability to detect mental workload across four distinct tasks: three well-validated cognitive psychology paradigms (Mental Rotation, Stroop, and N-Back) and a practical Chess [^1] We will hereafter capitalize “Chess” to refer to the task paradigm of playing chess puzzles. paradigm using chess puzzles. The selection of these four cognitive tasks: N-Back, Stroop, Mental Rotation, and Chess—creates a comprehensive framework for examining distinct yet interrelated aspects of cognitive function: the N-Back task primarily assesses working memory and information manipulation, the Stroop test measures inhibitory control and attention, the Mental Rotation task evaluates spatial reasoning, and the Chess task introduces a unique dimension of strategic thinking and pattern recognition that bridges and extends these cognitive domains. The Chess puzzle task presents a particularly valuable application in this battery because they require the simultaneous engagement of multiple cognitive processes: working memory (similar to N-Back, but in a more complex context), inhibitory control (as in Stroop, but applied to move selection), and spatial reasoning (complementing Mental Rotation, but with added strategic complexity).
@@
2 Background and related work
2.1 Low-Cost EEG for Mental Workload and State Classification
In recent years, advancements in consumer-grade, low-cost electroencephalography (EEG) devices—such as Muse II, OpenBCI, NeuroSky, InteraXon, and Emotiv—have positioned them as viable alternatives to traditional, expensive EEG systems (Vos et al. (2024)). These devices offer an accessible means of monitoring brain activity, enabling applications in dynamic, real-world settings. Their portability, affordability, and ease of use have made them practical tools for investigating mental workload and cognitive state classification, with demonstrated applications in attention monitoring, stress detection, and adaptive human-computer interaction (Knoll et al. (2011), Bashivan et al. (2016)).
Building upon this progress and aligning with our related work, studies such as (So et al. (2017)) have investigated the potential of single-channel EEG systems, such as NeuroSky, to differentiate nuanced cognitive tasks. Their evaluation of a single-probe frontal EEG system demonstrated its effectiveness in distinguishing tasks such as arithmetic operations, finger tapping, mental rotation, and lexical decisions. Notably, they observed an increase in $\theta$ band power during these tasks and successfully applied support vector machines (SVMs) to frequency-transformed 2.5-second windows, achieving classification accuracies exceeding 70% for arithmetic operations, finger tapping, and lexical decision. However, their cross-validation procedure did not account for auto-correlated features in time-related samples, highlighting opportunities for methodological refinement.
Devices like Muse II, which incorporate a four-channel dry electrode design, strike a balance between usability and functionality, making them well-suited for real-time applications. Despite this, these systems often face limitations in spatial resolution and signal fidelity when compared to research-grade devices (Lee et al. (2024)). To overcome such challenges, advancements in machine learning techniques—such as SVMs, convolutional neural networks (CNNs), and transformer-based models—have significantly improved classification accuracy and enhanced signal processing capabilities (Hassan et al. (2024),LaRocco et al. (2020), Beiramvand et al. (2023)).
@@ 2.1 Low-Cost EEG for Mental Workload and State Classification
2.2 Muse 2 in Research
Research has found that the MUSE 2 Headband, a portable and wireless device capable of measuring EEG, is a suitable alternative to traditional EEG methods. For example, to test the headband’s utility, a visual oddball paradigm experiment and a reward-learning task experiment were conducted. MUSE was able to observe and quantify the N200 and P300 ERP components in both tasks (Krigolson et al., 2017). The headband provides real-time feedback and monitoring of EEG indicators, and was validated for its accuracy in measuring EEG signals (Hawley et al., 2021; Krigolson et al., 2017).
Preliminary studies have also been done with the MUSE 2 which demonstrate the potential of low-cost EEG systems to classify mental states based on activity level. Bird et al. (2018), for instance, collected data from five participants across three mental states (relaxing, neutral, and concentrating), analyzing five different brainwave frequencies. Using feature selection techniques and machine learning classifiers including Bayesian Networks, Support Vector Machines (SVM) and Random Forests, they achieved over 87% accuracy in state classification. Another study done by Peining et al. (2017) investigated the development of a BCI system to enable drone control through consumer-grade EEG headsets. They focused on adapting drone thrust levels based on concentration, and likewise leveraged SVM and achieved 70% accuracy, however they also found wide variability among participants’ individual scores. In another study, Zhang and Cui (2022) tested the reliability of the MUSE 2 device against the TOBII Pro Nano eye tracking system. Their work suggests that the eye tracking system performs better, for classification was of N-Back levels 1 vs. 2. The relatively high cost of the TOBII device, however, likely puts it out of league with the MUSE 2.
@@ 2.3.1 N-Back
The conceptualization of the N-Back task traces its roots to the work of Kirchner in 1958 (Kirchner, 1958), establishing a cornerstone cognitive exemplar for the probing of working memory and executive function. This task is paramount in the realm of cognitive investigation due to its unique capacity to measure the retention and manipulation of information across brief intervals. Through the employment of diverse stimuli—ranging from auditory-verbal to visuospatial—and the increasing complexity introduced by 1-back, 2-back, and 3-back variations, the N-Back task presents an intricate measure of working memory capacity, executive function, and attentional control aptitudes. The operational framework of this task is meticulously designed to assess accuracy, response latencies, and error rates, thereby illuminating the rapidity and proficiency of cognitive operations by requiring participants to simultaneously store and manipulate information (Jaeggi et al., 2010; Conway et al., 2005; Kane and Engle, 2002).
2.3.2 Stroop
@@
As a widely used practice in psychology, the Stroop Test is a task designed to test the measurement of a user’s cognitive load (Gwizdka, 2010). Created by John Ridley Stroop in 1935, the task focuses on specific cognitive processes such as selective attention, memory, learning, cognitive load, and processing speed (MacLeod, 1991). The test requires the participant to identify the color of a printed word while ignoring the actual word itself. For instance, if the word ‘BLUE’ is printed in a yellow font, the correct response would be ‘yellow.’ The task measures the user’s accuracy and speed to give insight into the user’s cognitive ability and efficiency (MacLeod, 2015).
2.3.3 Mental Rotation
@@
The Mental Rotation Task (MRT) is a cognitive psychological test devised by Roger Shepard and Jacqueline Metzler in 1971, aimed at assessing spatial reasoning ability. In this test, participants are presented with two 3-dimensional objects in different perspectives positioned in space. The objective is to determine whether the objects are identical. Studies utilizing the MRT have revealed that reaction time tends to increase linearly as the angular disparity between the orientations grows, which indicates that participants mentally rotate one of the objects to determine if it matches the other (Shepard and Metzler, 1971). Moreover, EEG work has demonstrated that the MRT elicits activation in the superior parietal lobule and the intraparietal sulcus, indicating their involvement in spatial processing during the task (Osuagwu and Vuckovic, 2014).
2.3.4 Chess
@@
Several studies have explored the effects of chess playing on the brain. Increased theta power in posterior brain regions was observed when chess players engaged in faster-paced games, suggesting a link to long-term memory retrieval and chunk processing (Villafaina et al., 2019). High-level chess players have also been shown to exhibit increased alpha EEG power in the occipital area during chess playing, indicating a greater adaptive response to the cognitive demands of the task (Villafaina et al., 2021). Another study found that winning players exhibited higher theta power in frontal, central, and posterior brain regions as the difficulty of the opponent gradually increased; higher alpha power was observed in more challenging games, suggesting the engagement of creative thinking and the exploration of alternative solutions. By contrast, losing players demonstrated a decrease in beta and alpha power as the opponent’s difficulty escalated, indicating a lack of adaptive response to challenging situations and an inability to formulate effective solutions (Fuentes-García et al., 2019). Although one study explored ERP responses to the playing of Chess puzzles Hageman (2021), the application of a puzzle-paradigm in the context of EEG is nevertheless underexplored.
3 Materials and Methods
3.1 Study Outline
@@
3.2 Program Implementation
We implemented our experiments using JavaScript, and leveraged several key libraries and tools. We built the core experimental framework with jspsych (de Leeuw et al., 2023), which provided a flexible foundation for our study design. For the chess-related components, we employed chess.js (Schmich, 2023) to handle game logic and chessBoard.js (Oakman, 2023) to create an interactive visual interface. We interfaced our code with the MUSE 2 device by way of the muse-js (Urish, 2023) library.
3.3 Chess Puzzles Database
To create a database of Chess puzzles we began by downloading the lichess.org open chess puzzle database, which contains millions of Chess puzzles Lichess Team (2024a). To minimize confounding factors due to puzzle differences, we filtered the puzzles to only contain those that result in checkmate, which resulted in a total of 686,559 puzzles. Each puzzle had an associated Glicko2 rating to quantify its difficulty Glickman (2012). We further filtered the total puzzle set based on Glicko2 difficulty ratings within the range [600, 2250], where across the lichess.org site a rating of 600 represents the skill level of a complete beginner and the rating of 2250 represents the skill level of an expert (Lichess Team, 2024b). We then binned the puzzles into Glicko2 rating ranges of 50; the first bin contained puzzles with ratings from [600-650), and so on. The lower rating bins (600-1750) each contained over 10,000 puzzles. While the higher rating bins (1800-2250) contained fewer puzzles, each bin still maintained a minimum of 197 puzzles, ensuring sufficient availability for high-performing participants.
3.4 Task Details
3.4.1 Chess Task
@@
3.4.2 Cognitive Neuroscience Tasks
@@
Figure 2: The top row is a series of possible characters presented during a N-Back task. The bottom row represents the correct sequence if the task was a 2-Back, specifically. That is, on seeing all of the first four letters, the participant would press the ‘N’ key. However, the 5th character ‘G’ was also seen two characters prior, thus the participant would press ‘Y’. During the study, the characters are only shown one at a time, so the participant must remember N characters continuously.
@@
For the Mental Rotation task, participants were shown a series of pairs of 3D blocks. We used the block paradigm and images developed by (Ganis and Kievit, 2015). Participants were asked to press the y key if the blocks were the same but rotated, and the n key if they were not the same block. Blocks were the same with 50% probability. Each stimulus was shown for a maximum of 7500ms, after which the trial ended automatically. See Figure 4 for an example of the Rotation task.
Figure 4: Example of a trial used for the Mental Rotation task (images used from work of Ganis and Kievit (2015)). This is not a valid rotation, because the second object is not a rotated version of the first object.
@@
Figure 5: Experimental procedure per task. In the N-Back task, each set of 4 sub-blocks are a random permutation of 25 N-Back trials of the same N, where N is selected from {0,1,2,3}. Stroop blocks are 75 trials each, and Mental Rotation blocks are 24 trials each. The Chess blocks are each 30 puzzles in length. Note that this graphic only indicates the number of trials collected; due to data loss from the MUSE 2, time-per-trial, and number of trials per-block, our final dataset resulted in imbalanced samples (see Table 1).
3.5 Participant Information and Chess Player Skill

### v2

# Neural Signatures Within and Between Chess Puzzle Solving and Standard Cognitive Tasks for Brain-Computer Interfaces: A Low-Cost Electroencephalography Study

@@ Abstract
Consumer-grade electroencephalography (EEG) devices show promise for Brain-Computer Interface (BCI) applications, but their efficacy in detecting subtle cognitive states remains understudied. We developed a comprehensive study paradigm which incorporates a combination of established cognitive tasks (N-Back, Stroop, and Mental Rotation) and adds a novel ecological Chess puzzles task. We tested our paradigm with the MUSE 2, a low-cost consumer-grade EEG device. Using linear mixed-effects modeling we demonstrate successful distinctions of within-task workload levels and cross-task cognitive states based on the spectral power data derived from the MUSE 2 device. With machine learning we further show reliable predictive power to differentiate between workload levels in the N-Back task, and also achieve effective cross-task classification. These findings demonstrate that consumer-grade EEG devices like the MUSE 2 can be used to effectively differentiate between various levels of cognitive workload as well as among more nuanced task-based cognitive states, and that these tools can be leveraged for real-time adaptive BCI applications in practical settings.
© 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
I Introduction
@@
Electroencephalography (EEG) has emerged as a powerful tool in Brain-Computer Interface (BCI) research. In particular, EEG has an established capability to measure cognitive workload [1], which has proven valuable in ergonomics and human-computer interaction studies for differentiating task difficulties [2]. And recent advances in consumer-grade devices have expanded EEG’s accessibility beyond traditional clinical settings. This democratization has enabled researchers across diverse fields, from engineering to cognitive neuroscience, to leverage EEG technology in novel ways [3].
@@
However, bridging the gap between clinical neuroscience findings, initial validation of low-cost devices, and practical applications for general users remains an active area of research [4]. Further, while EEG technology continues to advance, significant work is still needed to establish reliable real-time signal processing methods suitable for BCI applications [5]. These challenges necessitate empirical studies to both validate the technological developments of lower-cost devices, and to inform the design of future user interfaces.
@@
Our research employs the MUSE 2, a wireless and portable EEG device initially validated for cognitive workload measurement in BCI applications [6, 7]. Despite this initial validation, there remains limited understanding of its practical application in realistic BCI scenarios. To address this gap, we investigate the device’s ability to detect mental workload across four distinct tasks: three well-validated cognitive psychology paradigms (Mental Rotation, Stroop, and N-Back) and a practical Chess [^1] We will hereafter capitalize “Chess” to refer to the task paradigm of playing chess puzzles. paradigm using chess puzzles. The selection of these four cognitive tasks: N-Back, Stroop, Mental Rotation, and Chess—creates a comprehensive framework for examining distinct yet interrelated aspects of cognitive function: the N-Back task primarily assesses working memory and information manipulation, the Stroop test measures inhibitory control and attention, the Mental Rotation task evaluates spatial reasoning, and the Chess task introduces a unique dimension of strategic thinking and pattern recognition that bridges and extends these cognitive domains. The Chess puzzle task presents a particularly valuable application in this battery because they require the simultaneous engagement of multiple cognitive processes: working memory (similar to N-Back, but in a more complex context), inhibitory control (as in Stroop, but applied to move selection), and spatial reasoning (complementing Mental Rotation, but with added strategic complexity).
@@
II Background and related work
II-A Low-Cost EEG for Mental Workload and State Classification
In recent years, advancements in consumer-grade, low-cost electroencephalography (EEG) devices—such as Muse II, OpenBCI, NeuroSky, InteraXon, and Emotiv—have positioned them as viable alternatives to traditional, expensive EEG systems [8]. These devices offer an accessible means of monitoring brain activity, enabling applications in dynamic, real-world settings. Their portability, affordability, and ease of use have made them practical tools for investigating mental workload and cognitive state classification, with demonstrated applications in attention monitoring, stress detection, and adaptive human-computer interaction [9, 10].
Building upon this progress and aligning with our related work, studies such as [11] have investigated the potential of single-channel EEG systems, such as NeuroSky, to differentiate nuanced cognitive tasks. Their evaluation of a single-probe frontal EEG system demonstrated its effectiveness in distinguishing tasks such as arithmetic operations, finger tapping, mental rotation, and lexical decisions. Notably, they observed an increase in $\theta$ band power during these tasks and successfully applied support vector machines (SVMs) to frequency-transformed 2.5-second windows, achieving classification accuracies exceeding 70% for arithmetic operations, finger tapping, and lexical decision. However, their cross-validation procedure did not account for auto-correlated features in time-related samples, highlighting opportunities for methodological refinement.
Devices like Muse II, which incorporate a four-channel dry electrode design, strike a balance between usability and functionality, making them well-suited for real-time applications. Despite this, these systems often face limitations in spatial resolution and signal fidelity when compared to research-grade devices [12]. To overcome such challenges, advancements in machine learning techniques—such as SVMs, convolutional neural networks (CNNs), and transformer-based models—have significantly improved classification accuracy and enhanced signal processing capabilities [13, 14, 15].
@@ II-A Low-Cost EEG for Mental Workload and State Classification
II-B Muse 2 in Research
Research has found that the MUSE 2 Headband, a portable and wireless device capable of measuring EEG, is a suitable alternative to traditional EEG methods. For example, to test the headband’s utility, a visual oddball paradigm experiment and a reward-learning task experiment were conducted. MUSE was able to observe and quantify the N200 and P300 ERP components in both tasks [16]. The headband provides real-time feedback and monitoring of EEG indicators, and was validated for its accuracy in measuring EEG signals [17, 16].
Preliminary studies have also been done with the MUSE 2 which demonstrate the potential of low-cost EEG systems to classify mental states based on activity level. [7], for instance, collected data from five participants across three mental states (relaxing, neutral, and concentrating), analyzing five different brainwave frequencies. Using feature selection techniques and machine learning classifiers including Bayesian Networks, Support Vector Machines (SVM) and Random Forests, they achieved over 87% accuracy in state classification. Another study done by [6] investigated the development of a BCI system to enable drone control through consumer-grade EEG headsets. They focused on adapting drone thrust levels based on concentration, and likewise leveraged SVM and achieved 70% accuracy, however they also found wide variability among participants’ individual scores. In another study, [18] tested the reliability of the MUSE 2 device against the TOBII Pro Nano eye tracking system. Their work suggests that the eye tracking system performs better, for classification was of N-Back levels 1 vs. 2. The relatively high cost of the TOBII device, however, likely puts it out of league with the MUSE 2.
@@ II-C1 N-Back
The conceptualization of the N-Back task traces its roots to the work of Kirchner in 1958 [19], establishing a cornerstone cognitive exemplar for the probing of working memory and executive function. This task is paramount in the realm of cognitive investigation due to its unique capacity to measure the retention and manipulation of information across brief intervals. Through the employment of diverse stimuli—ranging from auditory-verbal to visuospatial—and the increasing complexity introduced by 1-back, 2-back, and 3-back variations, the N-Back task presents an intricate measure of working memory capacity, executive function, and attentional control aptitudes. The operational framework of this task is meticulously designed to assess accuracy, response latencies, and error rates, thereby illuminating the rapidity and proficiency of cognitive operations by requiring participants to simultaneously store and manipulate information [20, 21, 22].
II-C2 Stroop
@@
As a widely used practice in psychology, the Stroop Test is a task designed to test the measurement of a user’s cognitive load [23]. Created by John Ridley Stroop in 1935, the task focuses on specific cognitive processes such as selective attention, memory, learning, cognitive load, and processing speed [24]. The test requires the participant to identify the color of a printed word while ignoring the actual word itself. For instance, if the word ‘BLUE’ is printed in a yellow font, the correct response would be ‘yellow.’ The task measures the user’s accuracy and speed to give insight into the user’s cognitive ability and efficiency [25].
II-C3 Mental Rotation
@@
The Mental Rotation Task (MRT) is a cognitive psychological test devised by Roger Shepard and Jacqueline Metzler in 1971, aimed at assessing spatial reasoning ability. In this test, participants are presented with two 3-dimensional objects in different perspectives positioned in space. The objective is to determine whether the objects are identical. Studies utilizing the MRT have revealed that reaction time tends to increase linearly as the angular disparity between the orientations grows, which indicates that participants mentally rotate one of the objects to determine if it matches the other [26]. Moreover, EEG work has demonstrated that the MRT elicits activation in the superior parietal lobule and the intraparietal sulcus, indicating their involvement in spatial processing during the task [27].
II-C4 Chess
@@
Several studies have explored the effects of chess playing on the brain. Increased theta power in posterior brain regions was observed when chess players engaged in faster-paced games, suggesting a link to long-term memory retrieval and chunk processing [28]. High-level chess players have also been shown to exhibit increased alpha EEG power in the occipital area during chess playing, indicating a greater adaptive response to the cognitive demands of the task [29]. Another study found that winning players exhibited higher theta power in frontal, central, and posterior brain regions as the difficulty of the opponent gradually increased; higher alpha power was observed in more challenging games, suggesting the engagement of creative thinking and the exploration of alternative solutions. By contrast, losing players demonstrated a decrease in beta and alpha power as the opponent’s difficulty escalated, indicating a lack of adaptive response to challenging situations and an inability to formulate effective solutions [30]. Although one study explored ERP responses to the playing of Chess puzzles [31], the application of a puzzle-paradigm in the context of EEG is nevertheless underexplored.
III Materials and Methods
III-A Study Outline
@@
III-B Program Implementation
We implemented our experiments using JavaScript, and leveraged several key libraries and tools. We built the core experimental framework with jspsych [32], which provided a flexible foundation for our study design. For the chess-related components, we employed chess.js [33] to handle game logic and chessBoard.js [34] to create an interactive visual interface. We interfaced our code with the MUSE 2 device by way of the muse-js [35] library.
III-C Chess Puzzles Database
To create a database of Chess puzzles we began by downloading the lichess.org open chess puzzle database, which contains millions of Chess puzzles [36]. To minimize confounding factors due to puzzle differences, we filtered the puzzles to only contain those that result in checkmate, which resulted in a total of 686,559 puzzles. Each puzzle had an associated Glicko2 rating to quantify its difficulty [37]. We further filtered the total puzzle set based on Glicko2 difficulty ratings within the range [600, 2250], where across the lichess.org site a rating of 600 represents the skill level of a complete beginner and the rating of 2250 represents the skill level of an expert [38]. We then binned the puzzles into Glicko2 rating ranges of 50; the first bin contained puzzles with ratings from [600-650), and so on. The lower rating bins (600-1750) each contained over 10,000 puzzles. While the higher rating bins (1800-2250) contained fewer puzzles, each bin still maintained a minimum of 197 puzzles, ensuring sufficient availability for high-performing participants.
III-D Task Details
III-D1 Chess Task
@@
III-D2 Cognitive Neuroscience Tasks
@@
B F G T G G G T X T L
Figure 2: A series of possible characters presented during a N-Back task. During the study, the characters are shown one at a time. Blue colored letters indicate 2-Back targets. That is, on seeing all of the first four letters, the participant would press the ‘N’ key. However, the 5th character ‘G’ was also seen two characters prior, thus the participant would press ‘Y’.
@@ B F G T G G G T X T L
For the Mental Rotation task, participants were shown a series of pairs of 3D blocks. We used the block paradigm and images developed by [39]. Participants were asked to press the y key if the blocks were the same but rotated, and the n key if they were not the same block. Blocks were the same with 50% probability. Each stimulus was shown for a maximum of 7500ms, after which the trial ended automatically. See Figure 4 for an example of the Rotation task.
Figure 4: Example of a trial used for the Mental Rotation task (images used from work of [39]). This is not a valid rotation, because the second object is not a rotated version of the first object.
@@
Figure 5: Experimental procedure per task. In the N-Back task, each set of 4 sub-blocks are a random permutation of 25 N-Back trials of the same N, where N is selected from {0,1,2,3}. Stroop blocks are 75 trials each, and Mental Rotation blocks are 24 trials each. The Chess blocks are each 30 puzzles in length. Note that this graphic only indicates the number of trials collected; due to data loss from the MUSE 2, time-per-trial, and number of trials per-block, our final dataset resulted in imbalanced samples (see LABEL:supp-tab:DATA_DESCRIPTION in the supplementary material).
III-E Participant Information and Chess Player Skill